1. Suppose that in an hour a German worker can produce 200 clothes or 20 cars, while a Japanese worker can produce 150 clothes or 10 cars. Each country has 8 hours available.       

(a) Suppose that without trade the labour in each country spend 25% of their time producing clothes. How many clothes and cars would Germany produce?  

2. Suppose that in an hour a German worker can produce 200 clothes or 20 cars, while a Japanese worker can produce 150 clothes or 10 cars. Each country has 8 hours available.       (b) Calculate the opportunity cost for cars in Japan and Germany.

3. Suppose that in an hour a German worker can produce 200 clothes or 20 cars, while a Japanese worker can produce 150 clothes or 10 cars. Each country has 8 hours available.      

(c) If these countries were open to trade, which country would export clothes? (1 mark) Explain.(2 marks)

(d) At what price should the two countries trade? (2 marks) Explain.(2 marks)
Answer #1

Answer)

As in an hour a german worker producee 200 clothes or 20 cars.

So, 25 % of 8 hours equal 2 hours.

In 2 hours, a german worker would produce:

Clothes = 200 × 2 = 400 units

Cars = 20 × 2 = 40 units

(b) Opportunity cost of car in germany :

= 200/20 clothes

= 10 clothes units

Oportunity cost of car in Japan:

= 150 / 10 clothes

= 15 clothes units

(c) To find the country which would exports the clothes, we need to find the opportunity cost of clothes in both the countries:

Germany :

  Opportunity cost of clothes:  

= 20/200  

= 0.1 cars

Japan:

Opportunity cost of clothes:

= 10/150  

= 0.07 cars

Japn would export clothes as it has the lower opportunity cost in the production of clothes.

(d)

The trade occurs within the range of the opportunity cost of a good in both the countries.

So, we can say that Germany and Japan would exchange clothes at a rate between the range - 0.07 and 0.1 cars per cloth.